A train starts from rest and moves with a constant acceleration of 2m/s^2 for half a minute. The brakes are then applied and the train comes to rest in one minute after applying brakes. Find (a) the total distance moved by the train, (b) the maximum speed attained by the train and (c) the position of the train at the half the maximum speed.

Grade:11

2 Answers

Himanshu Rathour
5 years ago
Let maximum speed attained by train is v, distance covered is s while accelerating and S while deaccelerating. 
v=0+2(30)
=60m/s
s=0+(1/2)(2)(900)
s=900m
0=60-a(60)
a=1m/s²
S=60(60)-½(60)²1
S=3600-1800
S=1800m
 
 
Himanshu Rathour
5 years ago
Sorry I misunderstood the c part so here is the answer to (c) part
Distance covered at half the maximum speed
30²=0+2(2)s1
s1=900/4
=225m